### v3

# Co-Engel graphs of certain finite non-Engel groups

Abstract.
Let $G$ be a group. Associate a graph $\mathcal{E}_{G}$ (called the co-Engel graph of $G$) with $G$ whose vertex set is $G$ and two distinct vertices $x$ and $y$ are adjacent if $x,{}_{k}y]\neq 1$ and $y,{}_{k}x]\neq 1$ for all positive integer $k$, where $x,{}_{k}y]$ is the iterated commutator $x,y,y,\ldots,y]$, with $k$ terms $y$ in the expression.
This graph, under the name “Engel graph”, was introduced by Abdollahi [2]. However, we argue that it is more naturally called the “co-Engel graph”.
Let $L(G)$ be the set of all left Engel elements of $G$. In this paper, we realize the induced subgraph of co-Engel graphs of certain finite non-Engel groups $G$ induced by $G\setminus L(G)$. We write $\mathop{\mathcal{E}_{c}^{-}}(G)$ to denote the subgraph of $\mathcal{E}_{G}$ induced by $G\setminus L(G)$. We also compute genus, various spectra, energies and Zagreb indices of $\mathop{\mathcal{E}_{c}^{-}}(G)$ for those groups. As a consequence, we determine (up to isomorphism) all finite non-Engel group $G$ such that the clique number $\omega(\mathop{\mathcal{E}_{c}^{-}}(G))$ is at most $4$ and $\mathop{\mathcal{E}_{c}^{-}}(G)$ is toroidal or projective. Further, we show that $\mathop{\mathcal{E}_{c}^{-}}(G)$ is super integral and satisfies the E-LE conjecture and the Hansen–Vukičević conjecture for the groups considered in this paper.
We also look briefly at the directed Engel graph, with an arc $x\to y$ if $y,{}_{k}x]=1$ for some $k$. We show that, if $G$ is a finite soluble group, this graph either is the complete directed graph (which occurs only if $G$ is nilpotent), or has pairs of vertices joined only by single arcs. We also show that the (directed or undirected) Engel graph of a group $G$ is the lexicographic product of a complete graph of order $Z_{\infty}(G)$ by the (directed or undirected) Engel graph of $G/Z_{\infty}(G)$, where $Z_{\infty}(G)$ is the hypercenter of $G$.
Key words and phrases:
Engel graph; Finite group
2010 Mathematics Subject Classification:
Primary 20D60; Secondary 05C25
*Corresponding author
1. Introduction
The term “graphs defined on groups” refers to graphs whose vertex set is the group $G$ (or a suitable subset) with the adjacency of elements $x$ and $y$ defined in terms of group-theoretic properties of these elements. The first such graph was the commuting graph of a group, with $x$ and $y$ joined if $xy=yx$. This was used by Brauer and Fowler in a 1955 paper [7] containing a foundational result for the study of finite simple groups (they used the ideas of graph theory but not the name). The graphs defined here can be regarded as a generalisation of the commuting graph. In the last few decades, many more such graphs have been defined, including the power graph and enhanced power graph, nilpotency graph, generating graph, independence graph, and rank graph. Our topic here is the Engel graph, or more precisely its complement.
In a finite group $G$, the commutator of two elements $x,y$ is given by
$x,y]=x^{-1}y^{-1}xy,$
and for a positive integer $k$, the $k$th Engel element is defined inductively by
$x,{}_{1}y]=[x,y],\qquad[x,{}_{k}y]=[[x,{}_{k-1}y],y].$
We say that $(x,y)$ satisfies the Engel condition if $x,{}_{k}y]=1$ for some $k$.
Abdollahi [2] defined the “Engel graph” of a group $G$, with $x,y$ joined if they don’t satisfy the Engel condition in either order, and the left Engel elements (isolated in this graph) deleted. He showed that for infinite soluble non-Engel groups, this graph has no infinite clique if and only if there is a finite bound on its clique number (this is the analogue of a similar result by Neumann [16] for the non-commuting graph, answering a question of Erdős, and depends on a result of Longobardi and Maj [15]). He also showed that a finite group is soluble if the graph has clique number at most $15$; moreover, the index of the Fitting subgroup is bounded by a function of the clique number. Finally, he studied the connectedness of the graph.
The study languished for some time until it was revived in the present decade. Recent authors have changed the definition so that Abdollahi’s graph is the complement of the Engel graph as now defined. (Abdollahi gave the first author of this paper his blessing for this change of usage in [8].) Also, unlike most graphs defined on groups (the enhanced power graph, commuting graph, nilpotent graph, generating graph, independence graph, and others), but similar to the situation with the power graph, the Engel graph arises naturally as a directed graph, with an arc $x\to y$ if $x,{}_{k}y]=1$ for some integer $k$. The undirected version is obtained by ignoring directions (and regarding two oppositely directed arcs between the same pair of vertices as a single undirected edge).
Detomi, Lucchini and Nemmi [10] showed that the Engel digraph of a finite group $G$ is weakly connected (with undirected diameter at most $10$), and is strongly corrected if $G/Z_{\infty}(G)$ is neither almost simple nor a Frobenius group. Moreover, if G is soluble and $G/Z_{\infty}(G)$ is not Frobenius, the diameter is at most $4$. Della Volta, Mastrogiacomo and Spiga [9] consider the case where $G$ is almost simple and show that, in most cases, strong connectedness holds here too. Finally, the first author [8] used the Engel graph to show that the set of dominating vertices in the nilpotency graph of a finite group $G$ is the hypercentre $Z_{\infty}(G)$.
In this paper, we return to the study of Abdollahi’s graph, which we re-name as the co-Engel graph. We compute various properties of this graph including genus, spectra, energies and Zagreb indices for a selection of groups, showing that the graph is super integral and satisfies the E-LE conjecture and the Hansen–Vukičević conjecture.
After a section giving definitions and assumed results, we show that the conditions on a finite group that its nilpotency graph is complete, its Engel graph is complete, and these two graphs are equal are all equivalent to the statement that the group is nilpotent. We also show that the Engel graph of a group with hypercenter $Z_{\infty}(G)$ is the lexicographic product of a complete graph of order $|Z_{\infty}(G)|$ by the Engel graph of $G/Z_{\infty}(G)$.
In other words, the co-Engel graph is null preciselly for nilpotent groups, so we exclude these from our discussion and remove the isolated vertices of the co-Engel graph in order to study its properties.
2. Definitions and basic results
We mostly use basic terminology from group theory and graph theory, giving references where necessary. Note that we denote an arc from $x$ to $y$ in a directed graph by the standard graph-theoretic notation $x\to y$, not $x\mapsto y$ as used in some recent papers on the Engel graph.
The directed Engel graph $\mathop{\overrightarrow{\mathcal{E}}}(G)$ of $G$ has vertex set $G$, with an arc $x\to y$ if $y,{}_{n}x]=1$ for some natural number $n$. The undirected Engel graph $\mathop{\mathcal{E}}(G)$ has an edge $\{x,y\}$ if either $x\to y$ or $y\to x$ (or both) in the directed Engel graph. The full co-Engel graph $\mathop{\mathcal{E}_{c}}(G)$ is the complement of the undirected Engel graph: that is, it has an edge $\{x,y\}$ if neither $x\to y$ nor $y\to x$ in the directed Engel graph.
Baer [3] showed that, in our terminology, the set of elements $x\in G$ with the property that $x\to y$ for all $y\in G$ with $y\neq x$ is the Fitting subgroup of $G$, the largest nilpotent normal subgroup of $G$. Abdollahi [2] showed a stronger result:
Proposition 2.1.
The set of vertices $x$ which are joined to all others in the Engel graph of $G$ is the Fitting subgroup of $G$.
Often we delete the isolated vertices and refer to the result as the co-Engel graph, denoted $\mathop{\mathcal{E}_{c}^{-}}(G)$.
The nilpotency graph $\mathop{\mathcal{N}}(G)$ of $G$ has an edge $\{x,y\}$ if the group $\langle x,y\rangle$ generated by these two elements is nilpotent.
A nilpotent group satisfies the commutator identity $x_{1},x_{2},\ldots,x_{k}]=1$ for all $x,y$ (for some natural number $k$), where the iterated commutator is defined by a similar induction
$x_{1},x_{2},\ldots,x_{k}]=[[x_{1},x_{2},\ldots,x_{k-1}],x_{k}]$
for $k\geq 3$; the smallest such $k$ is its nilpotency class. Thus we see that any nilpotent group is an Engel group. An important theorem of Zorn [25] states that the converse holds for finite groups. It follows that the nilpotency graph of a finite group $G$ is complete if and only if its directed Engel graph is the complete digraph. We will strengthen this result for finite soluble groups in the next section.
Zorn’s result shows that the terms “non-nilpotent group” and “non-Engel group” are synonymous for finite groups, and we will use them interchangeably.
The hypercenter of $G$ is defined as follows. The lower central series of $G$ is the sequence of subgroups
$\{e\}=Z_{0}(G)\leq Z_{1}(G)\leq Z_{2}(G)\leq\cdots,$
where $Z_{i+1}(G)/Z_{i}(G)$ is the center of $G/Z_{i}(G)$. The hypercenter is the union of the subgroups of this series. In a finite group, the series terminates, and $Z_{\infty}(G)=Z_{n}(G)$ for some $n$. Thus, the hypercenter of a finite group is nilpotent (and normal in $G$), and so is contained in the Fitting subgroup of $G$ (the maximal normal nilpotent subgroup of $G$).
3. Properties of the Engel (di)graph
We have two purposes in this section: we show that the Engel (di)graph of $G$ is complete if and only if $G$ is nilpotent; and we give the structure of the Engel (di)graph in terms of the Engel (di)graph of $G/Z_{\infty}(G)$.
We recall that the directed Engel graph of $G$ has vertex set $G$, with an arc $x\to y$ if $y,{}_{n}x]=1$ for some natural number $n$. We show that, if it is not the complete directed graph, then it has pairs of vertices which are joined by a single arc in this digraph. We note that there are three types of pairs $\{x,y\}$: joined by arcs in both directions; joined by an arc in one direction but not the other; or not joined by an arc (the last type are the edges of the co-Engel graph. We will see that, if $x,{}_{n}y]=[y,{}_{m}x]=1$, there is no bound for $m$ in terms of $n$.
By definition, the set $L(G)$ of left Engel elements in $G$ (the isolated vertices in $\mathop{\mathcal{E}_{c}^{-}}(G)$) can be described as the set
$\{x\in G:(\forall y\in G)(x\neq y\Rightarrow x\to y)\}$
of vertices which dominate all other vertices in the directed Engel graph. If $G$ is a finite group, then $L(G)$ is the Fitting subgroup of $G$ (the maximal normal nilpotent subgroup), see Baer [3]. In particular, if $G$ is nilpotent, then its directed Engel graph is the complete directed graph, with arcs in both directions between any pair of vertices. Note that every edge $\{x,y\}$ of the nilpotency graph has arcs in both directions in $\mathop{\overrightarrow{\mathcal{E}}}(G)$, and so is an edge of the Engel graph. However, if there is a single arc $x\to y$ from $x$ to $y$. then $y,{}_{k}x]\neq 1$ for all integers $k$; thus $\langle x,y\rangle$ is not nilpotent, and so $\{x,y\}$ is an edge of the Engel graph but not the nilpotency graph.
We now show that, except in the case where $G$ is nilpotent, there exist single arcs in the directed Engel graph.
Lemma 3.1.
Let $G$ be a finite soluble group. If the directed Engel graph of $G$ has no single arcs, then $G$ is nilpotent (and the graph is complete).
Proof.
Let $F$ be the Fitting subgroup of $G$. Then $F$ is the set of left Engel elements, so we have an arc $x\to y$ for all $x\in F$, $y\in G$ with $x\neq y$. If there are no single edges, then we also have an arc $y\to x$ for all such $x,y$.
If $F=G$, we are done; so suppose not. Let $H$ be a subgroup of $G$ such that $H/F$ is a minimal normal subgroup of $G/F$. Then $H/F$ is an elementary abelian $p$-group, for some prime $p$. We will show that $H$ is nilpotent, contradicting the fact that $F$ is the maximal normal nilpotent subgroup of $G$.
Take an element $x\in H\setminus F$. We can assume that the order of $x$ is a power of $p$. Let $Q$ be a non-trivial Sylow $q$-subgroup of $F$, where $q\neq p$. If $x$ acts non-trivially on $Q$, then it acts non-trivially on $Q/\Phi(Q)$, by the Burnside basis theorem [13, Chapter 10], where $\Phi(Q)$ is the Frattini subgroup of $Q$. This quotient is an elementary abelian $q$-group, and is completely reducible as an $\langle x\rangle$-module, since $q\neq p$.
Choose an element $y\in Q$ such that $y\Phi(Q)$ lies in a submodule $M$ of $Q/\Phi(Q)$ on which $x$ acts non-trivially. Then $x^{-1}yx\Phi(Q)\in M$ and $x^{-1}yx\Phi(Q)\neq y\Phi(Q)$, so $y,x]\Phi(Q)$ is a non-zero element of $M$. Now the same argument applies, with induction, to show that $y,{}_{n}x]\Phi(Q)\neq\Phi(Q)$, and hence $y,{}_{n}x]\neq 1$, for all $n$. But this contradicts the fact that $x\to y$ in the directed Engel graph.
This shows that a Sylow $p$-subgroup $P$ of $H$ acts trivially on each Sylow $q$-subgroup $Q$ of $F$ (and hence of $H$) for $q\neq p$ (for elements of $P\cap F$ commute with $Q$ since $F$ is nilpotent, and the remaining elements by the argument we have just given). Since $F$ is nilpotent, its distinct Sylow subgroups commute with each other; so all the distinct Sylow subgroups of $H$ commute with each other, and $H$ is nilpotent.
This contradiction completes the proof. ∎
Note that the single arcs guaranteed by this theorem have the form $x\to y$ where $x\in L(G)$ and $y\in G\setminus L(G)$.
For the next result, we recall that the nilpotency graph of the finite group $G$ has an edge $\{x,y\}$ if and only if $\langle x.y\rangle$ is nilpotent.
Theorem 3.2.
Let $G$ be a finite soluble group. Then the following are equivalent:
  1. (a)
    $G$ is nilpotent;
  2. (b)
    the nilpotency graph $\mathop{\mathcal{N}}(G)$ is complete;
  3. (c)
    the Engel graph $\mathop{\mathcal{E}}(G)$ is complete;
  4. (d)
    the nilpotency and Engel graphs of $G$ are equal.
Proof.
If $G$ is nilpotent, then its $2$-generator subgroups are nilpotent and hence Engel, so (b) – (d) all hold. We prove the converses.
Suppose that $G$ is not nilpotent. By Lemma 3.1, the Engel digraph has two vertices $x$ and $y$ joined by a single arc. Then one of $x,{}_{k}y]=1$ and $y,{}_{k}x]=1$ fails for all $k$. Thus, $\langle x,y\rangle$ is not Engel, and hence not nilpotent, so the nilpotency graph is not complete. Thus (b) implies (a). Moreover, $\{x,y\}$ is an edge in the Engel graph but not in the nilpotency graph; so also (d) implies (a).
Finally, suppose that $G$ is not nilpotent. The minimal non-nilpotent groups were classified by Schmidt [18]; a convenient description can be found in [4]. Each of them involves a cyclic $q$-group $Q$ acting fixed-point-freely on an abelian $p$-group $P$. Let $x$ be a generator of $Q$, and $y$ a conjugate of $x$ by an element of $P$. Then $[x,y]$ is a non-zero element of $P$, and so $x,{}_{k}y]$ is a non-zero element of $P$ for all $k$. Similarly with $x$ and $y$ interchanged. So $x$ and $y$ are not joined in the Engel graph of $G$. Thus (c) implies (a). ∎
Note that the single arcs in the above proof join a vertex in $F(G)$ to a vertex outside $F(G)$. We will see later that, if $G$ is a dihedral group, then every single arc is of this form. However, this is not always the case: there can be single arcs outside $L(G)$. Consider the group $G=S_{4}$. In this case, $L(G)$ is the Fitting subgroup of $G$, consisting of the identity and the three double transpositions. Now $G$ has a subgroup isomorphic to $S_{3}$, all of whose non-identity elements lie outside $\mathop{\mathrm{Fit}}(G)$, and the induced subgraph on these vertices contains single arcs(from $x$ to $y$ if $x$ and $y$ have orders $3$ and $2$ respectively).
Problem 3.1.
Suppose that two groups $G$ and $H$ have isomorphic Engel graphs. Do they necessarily have isomorphic directed Engel graphs?
The main theorem of this section shows that this is true if the Engel graph is complete.
Now we turn to the connection of the Engel graphs of a group $G$ and of $G/Z_{\infty}(G)$. The lexicographic product of graphs $\Gamma_{1}$ and $\Gamma_{2}$ is the graph whose vertex set is $v(\Gamma_{1})\times v(\Gamma_{2})$, which we regard as a disjoint union of copies $C_{x}$ of $v(\Gamma_{1})$ indexed by $\Gamma_{2}$. We place a copy of $\Gamma_{1}$ on each of these sets. For $x\neq y$ if there is an edge (resp. arc) from $x$ to $y$ in $\Gamma_{2}$, then we put all edges (resp. arcs) from $C_{x}$ to $C_{y}$; otherwise we put no edges.
Theorem 3.3.
Let $G$ be a finite group.
  1. (a)
    The Engel digraph of $G$ is isomorphic to the lexicographic product of a complete directed graph of order $|Z_{\infty}(G)|$ by the Engel digraph of $G/Z_{\infty}(G)$.
  2. (b)
    The Engel graph of $G$ is isomorphic to the lexicographic product of a complete graph of order $|Z_{\infty}(G)|$ by the Engel graph of $G/Z_{\infty}(G)$.
  3. (c)
    The Engel digraph of $G$ is isomorphic to the lexicographic product of a null graph of order $|Z_{\infty}(G)|$ by the Engel digraph of $G/Z_{\infty}(G)$.
Proof.
Parts (b) and (c) follow immediately from (a), so it suffices to prove (a). We prove this by induction on the length of the lower central series of $G$.
Let $Z=Z(G)$ be the center of $G$. The elements of $G/Z$ are the cosets of $Z$ in $G$. If $x,{}_{k}y]=1$, then $xZ,{}_{k}yZ]=Z$. Conversely, if $xZ,{}_{k}yZ]=Z$ in $G/Z$, then $x,{}_{k}y]\in Z$ in $G$, and so $x,{}_{k+1}y]=1$. Moreover, the induced subgraph on $Z$ is complete. This demonstrates the result in this case.
Now for $G/Z_{n}(G)$ with $n>1$, the result follows easily, using the result for $G/Z_{n-1}(G)$, since we have $G/Z_{n}(G)\cong(G/Z_{n-1}(G))/(Z_{n}(G)/Z_{n-1}(G))$. So the theorem is proved. ∎
We will use this theorem a number of times in what follows. It will also be useful in further research: it implies that, in studying groups $G$ whose co-Engel graph is perfect, or chordal, or a cograph, or in investigating the effect of twin reduction, we may assume without loss that the center of $G$ is trivial.
Other results about the Engel digraph do not extend to the undirected case. For example, the strong product of digraphs $D_{1}$ and $D_{2}$ has vertex set $v(D_{1})\times v(D_{1})$, with an arc $(a,x)\to(b,y)$ if either $a=b$ or $a\to b$, and either $x=y$ or $x\to y$, but not both $a=b$ and $x=y$.
Proposition 3.4.
Let $G$ and $H$ be finite groups. Then the Engel digraph of $G\times H$ is the strong product of the Engel digraphs of $G$ and $H$.
The proof is straightforward. However, it does not hold for the Engel and co-Engel graphs. For, if there are single arcs $a\to b$ and $y\to x$, then the strong product of the Engel graphs has an edge $\{(a,x),(b,y)\}$ which is not an edge of the Engel graph of $G\times H$. (This problem does not arise if one of the factors is an Engel group, since its Engel digraph has no single arcs. Indeed, if $G$ is nilpotent, then it is contained in the hypercenter of $G\times H$.)
4. Realization of $\mathop{\mathcal{E}_{c}^{-}}(G)$
From this point on, we turn our attention to properties of co-Engel graphs. In this section we realize certain graphs as $\mathop{\mathcal{E}_{c}^{-}}(G)$ for certain finite groups. Let $K_{n}$ be the complete graph on $n$ vertices and $K_{n_{1},n_{2},\dots,n_{m}}$ be the complete $m$-partite graph with parts of size $n_{1},n_{2},\dots,n_{m}$. We write $K_{m\cdot n}=K_{n_{1},n_{2},\dots,n_{m}}$ if $n_{1}=n_{2}=\cdots=n_{m}=n$.
It was shown in [2] that $\mathop{\mathcal{E}_{c}^{-}}(D_{6})\cong K_{3}$ and $\mathop{\mathcal{E}_{c}^{-}}(D_{12})\cong K_{3\cdot 2}$. In the following theorem we shall generalize these results and realize $\mathop{\mathcal{E}_{c}^{-}}(G)$, with $G$ is a dihedral group $D_{2m}$ or $D_{2^{t+1}m}$, or a generalised quaternion group $Q_{2^{t+1}m}$, where $t\geq 1$ and $m\geq 3$ is odd.
Theorem 4.1.
If $t\geq 1$ and $m\geq 3$ is odd, then $\mathop{\mathcal{E}_{c}^{-}}(D_{2^{t+1}m})$ and $\mathop{\mathcal{E}_{c}^{-}}(Q_{2^{t+1}m})$ are isomorphic to $K_{m\cdot 2^{t}}$. Further, $\mathop{\mathcal{E}_{c}^{-}}(D_{2m})$ is isomorphic to $K_{m}$.
Proof.
It is not difficult to see that if $G$ denotes one of the groups $D_{2^{t+1}m}$ and $Q_{2^{t+1}m}$, then $L(G)=\mathop{\mathrm{Fit}}(G)=\langle y\rangle$, where $\mathop{\mathrm{Fit}}(G)$ is the Fitting subgroup of $G$ (see [3]). Thus $G\setminus L(G)=\{x,xy,\dots,xy^{2^{t}m-1}\}$. We have
[…]
Thus, $xy^{i},{}_{n}xy^{j}]=1$ if and only if $2^{t}m\mid 2^{n}(i-j)$, that is $i\equiv j\mod m$. Let
$A_{j}=\{xy^{i}\mid 0\leq i\leq 2^{t}m-1,i\equiv j\mod m\}$
for $j=0,1,\ldots,m-1$. Then $A_{0},A_{1},\dots,A_{m-1}$ are all disjoint subsets of $G\setminus L(G)$ such that $|A_{i}|=2^{t}$ for all $0\leq i\leq m-1$. Notice that $u\in A_{i}$ and $v\in A_{j}$ are adjacent if and only if $i\neq j$. Therefore, $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{2^{t},\dots,2^{t}}=K_{m\cdot 2^{t}}$.
On the other hand, for $D_{2m}$ (where $m$ is odd) we have $|A_{i}|=1$ for all $0\leq i\leq m-1$. Hence, $\mathop{\mathcal{E}_{c}^{-}}(D_{2m})\cong K_{m}$. ∎
In the next theorem, we realize $\mathop{\mathcal{E}_{c}^{-}}(G)$ when $G=F_{p,q}$, the non-abelian group of order $pq$ where $p$ and $q$ are primes.
Theorem 4.2.
If $G$ is the group $F_{p,q}=\langle a,b\mid a^{p}=b^{q}=1,a^{-1}ba=b^{r}\rangle$ of order $pq$, where $p$ and $q$ are primes such that $q\equiv 1\mod p$ and $r^{p}\equiv 1\mod q$, then $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{q\cdot(p-1)}$.
Proof.
We have  $F_{p,q}=\{1,b,b^{2},\dots,b^{q-1},a,a^{2},\dots,a^{p-1},ab,\dots,$ $ab^{q-1},\dots,a^{p-1}b,\dots,$ $a^{p-1}b^{q-1}\}$. Thus, if $g\in F_{p,q}$ then $g=a^{i}b^{j}$, where $0\leq i\leq p-1$ and $0\leq j\leq q-1$.
We also have $a^{i}b^{t}a^{j}b^{s}=a^{i+j}b^{s+tr^{j}}$. Therefore, if $i\neq 0$ then $(a^{i}b^{j})^{p}=1$. Note that the group $F_{p,q}$ is the union of its subgroups $H_{b}=\langle b\rangle=\mathop{\mathrm{Fit}}(F_{p,q})=L(F_{p,q})$ and
$H_{i}=\langle ab^{i}\rangle=\{1,ab^{i},a^{2}b^{i(r+1)},a^{3}b^{i(r^{2}+r+1)},\dots,a^{p-1}b^{i(r^{(p-1)-1}+r^{(p-1)-2}+\dots+r+1)}\},$
where $0\leq i\leq q-1$. It is easy to see that the intersection of any two of these subgroups is $Z(F_{p,q})=\{1\}$.
Let $x=a^{t}b^{i(r^{t-1}+\dots+r+1)}=a^{t}b^{it^{\prime}}\in H_{i}$ and $y=a^{s}b^{j(r^{s-1}+\dots+r+1)}=a^{t}b^{js^{\prime}}\in H_{j}$, where $t^{\prime}=r^{t-1}+\dots+r+1$ and $s^{\prime}=r^{s-1}+\dots+r+1$. If $i=j$, then $[x,y]=1$. So suppose that $i\neq j$. We have
$x,y]=b^{it^{\prime}(r^{s}-1)+js^{\prime}(1-r^{t})}=b^{k},$
where $k={it^{\prime}(r^{s}-1)+js^{\prime}(1-r^{t})}$. Thus
$x,{}_{2}y]=[b^{k},a^{s}b^{js^{\prime}}]=b^{k(r^{s}-1)}=b^{l},$
where $l=k(r^{s}-1)$. Therefore.
$x,{}_{3}y]=[b^{l},a^{s}b^{js^{\prime}}]=b^{l(r^{s}-1)}=b^{k(r^{s}-1)^{2}}.$
In general, we have $x,{}_{n}y]=b^{k(r^{s}-1)^{n-1}}$.
Again,
$k$ $=it^{\prime}(r^{s}-1)+js^{\prime}(1-r^{t})$
$=i(t^{\prime}r^{s}-t^{\prime})+j(s^{\prime}-r^{t}s^{\prime})=i(t^{\prime}r^{s}+s^{\prime}-s^{\prime}-t^{\prime})+j(s^{\prime}-r^{t}s^{\prime}-t^{\prime}+t^{\prime})$
$=i((t+s)^{\prime}-s^{\prime}-t^{\prime})+j(s^{\prime}-(t+s)^{\prime}+t^{\prime}),$
where  $(t+s)^{\prime}=r^{t+s-1}+\dots+r^{s+1}+r^{s}+r^{s-1}+\dots+r+1$. Thus
$k=(i-j)((t+s)^{\prime}-s^{\prime}-t^{\prime})$ $=(i-j)\left(\frac{r^{t+s}-1}{r-1}-\frac{r^{s}-1}{r-1}-\frac{r^{t}-1}{r-1}\right)$
$=\frac{(i-j)(r^{t}-1)(r^{s}-1)}{r-1}.$
Clearly $q$ does not divide $(i-j),(r^{t}-1)$ and $(r^{s}-1)$, since $1\leq i,j\leq q-1$, $i\neq j$ and $0\leq r,s\leq p-1$. Therefore, $q\nmid k(1-r^{s})^{n-1}$ and so $x,{}_{n}y]\neq 1$ for any integer $n$. Hence, the result follows. ∎
The following remark is useful in realizing $\mathop{\mathcal{E}_{c}^{-}}(G)$ if $G$ is a direct product of two groups.
Remark 4.1.
  1. (a)
    Let $G$ and $H$ be two groups and $(u,x),(v,y)\in G\times H$. Then
    $(u,x),(v,y)]=(u^{-1}v^{-1}uv,x^{-1}y^{-1}xy)=([u,v],[x,y])$
    and so
[…]
    $=[([u,v],[x,y]),(v,y)]$
    $=([[u,v],v],[[x,y],y])=([u,{}_{2}v],[x,{}_{2}y]).$
    In general, we have
    $(u,x),{}_{n}[v,y])=([u,{}_{n}v],[x,{}_{n}y]).$
  2. (b)
    Let $G$ be a group and $x,y\in G$. If $x,{}_{n}y]=1$ for some positive integer $n$, then $x,{}_{n+1}y]=[[x,{}_{n}y],y]=[1,y]=1$. Thus $x,{}_{m}y]=1$ for all $m\geq n$.
Theorem 4.3.
Let $G$ be a finite non-Engel group such that $\mathop{\mathcal{E}_{c}^{-}}(G))\cong K_{m\cdot n}$. Let $H$ be a finite Engel group with $|H|=l$. Then $\mathop{\mathcal{E}_{c}^{-}}(H\times G)\cong K_{lm\cdot n}$.
Proof.
Let $(u,x),(v,y)\in H\times G$. Then, by the Remark 4.1(a), we have
[…]
Since $H$ is a finite Engel group, there exist positive integers $i$ and $j$, such that $u,{}_{i}v]=1$ and $v,{}_{j}u]=1$. Suppose $x,{}_{a}y]=1$ and $y,{}_{b}x]=1$ for some positive integers $a$ and $b$. Let $r\geq\max\{i,a\}$ and $s\geq\max\{j,b\}$. Then, by Remark 4.1(b), we have $([u,{}_{r}v],[x,{}_{r}y])=1$ and $([v,{}_{s}u],[y,{}_{s}x])=1$. Hence, it follows that $([u,{}_{k}v],[x,{}_{k}y])\neq 1$ and $([v,{}_{k}u],[y,{}_{k}x])\neq 1$ if and only if $x,{}_{k}y]\neq 1$ and $y,{}_{k}x]\neq 1$ for all positive integer $k$. Thus, $L(H\times G)=H\times L(G)$ and hence $v(\mathop{\mathcal{E}_{c}^{-}}(H\times G))=H\times(G\setminus L(G))$.
Let $G_{1},G_{2},\dots,G_{m}$ be the partite sets of $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m\cdot n}$. Since $([u,{}_{k}v],[x,{}_{k}y])$ $\neq 1$ and $([v,{}_{k}u],[y,{}_{k}x])\neq 1$ if and only if $x,{}_{k}y]\neq 1$ and $y,{}_{k}x]\neq 1$ for all positive integer $k$, we have $\mathop{\mathcal{E}_{c}^{-}}(H\times G)$ is isomorphic to $K_{lm\cdot n}$ with partite sets $H\times G_{1},H\times G_{2},\dots,H\times G_{m}$. This completes the proof. ∎
5. Genus of $\mathop{\mathcal{E}_{c}^{-}}(G)$
In this section, we determine the genus of $\mathop{\mathcal{E}_{c}^{-}}(G)$ for the groups considered in Section 4. It is well-known that
$K_{n}=\left\lceil\frac{(n-3)(n-4)}{12}\right\rceil$ (1)
for $n\geq 3$. If $K_{m,n}$ denotes the complete bipartite graph with parts of size $m,n\geq 2$, then
$K_{m,n}=\left\lceil\frac{(m-2)(n-2)}{4}\right\rceil.$ (2)
From the main theorem of [23], we also have
$K_{mn,n,n}=\frac{(mn-2)(n-1)}{2}$ (3)
for all positive integers $m$ and $n$. The following is useful in our computation.
Theorem 5.1.
If $\Gamma=K_{a\cdot b}$, where $a\geq 3,b\geq 2$, then $\gamma(\Gamma)=\frac{a(a-1)}{2}\left\lceil\frac{(b-2)^{2}}{4}\right\rceil+\left\lceil\frac{(a-3)(a-4)}{12}\right\rceil$.
Proof.
Follows from Theorem 4.1 of [5]. ∎
Theorem 5.2.
If $t\geq 1,m\geq 3$, and $m$ is odd, then the genus of $\mathop{\mathcal{E}_{c}^{-}}(G)$, where $G=D_{2^{t+1}m}$ and $Q_{2^{t+1}m}$, are given by
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=\frac{m(m-1)(2^{t-1}-1)^{2}}{2}+\left\lceil\frac{(m-3)(m-4)}{12}\right\rceil.$
Further, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(D_{2m}))=\left\lceil\frac{(m-3)(m-4)}{12}\right\rceil$.
Proof.
By Theorem 4.1 we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m\cdot 2^{t}}$, where $G=D_{2^{t+1}m}$ and $Q_{2^{t+1}m}$. Therefore, putting $a=m$ and $b=2^{t}$ in Theorem 5.1, we get
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=\frac{m(m-1)}{2}\left\lceil\frac{(2^{t}-2)^{2}}{4}\right\rceil+\left\lceil\frac{(m-3)(m-4)}{12}\right\rceil.$
Hence, the first part of the result follows. The second part follows from (1) using the fact that $\gamma(\mathop{\mathcal{E}_{c}^{-}}(D_{2m}))=\gamma(K_{m})$. ∎
Theorem 5.3.
If $F_{p,q}=\langle a,b\mid a^{p}=b^{q}=1,a^{-1}ba=b^{r}\rangle$, where $p$ and $q$ are primes, $q\equiv 1\mod p$ and $r^{p}\equiv 1\mod q$ then
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(F_{p,q}))=\begin{cases}\left\lceil\frac{(q-3)(q-4)}{12}\right\rceil,&\text{ if }p=2\vskip 5.69046pt\\ \frac{q(q-1)}{2}\left\lceil\frac{(p-3)^{2}}{4}\right\rceil+\left\lceil\frac{(q-3)(q-4)}{12}\right\rceil,&\text{ if }p\geq 3.\end{cases}$
Proof.
By Theorem 4.2 we have $\mathop{\mathcal{E}_{c}^{-}}(F_{p,q})\cong K_{q\cdot(p-1)}$. If $p=2$ then $\mathop{\mathcal{E}_{c}^{-}}(F_{p,q})\cong K_{q}$ noting that $K_{q\cdot 1}=K_{q}$. Therefore, the result follows from (1).
If $p\geq 3$ then, putting $a=q$ and b=p1b=p-1 in Theorem 5.1, we get
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(F_{p,q}))=\frac{q(q-1)}{2}\left\lceil\frac{((p-1)-2)^{2}}{4}\right\rceil+\left\lceil\frac{(q-3)(q-4)}{12}\right\rceil.$
Hence, the result follows. ∎
Theorem 5.4.
Let $G$ be a finite non-Engel group such that $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m\cdot n}$. Let $H$ be a finite Engel group with $|H|=l$ and $X=H\times G$. Then
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(X))=\frac{lm(lm-1)}{2}\left\lceil\frac{(n-2)^{2}}{4}\right\rceil+\left\lceil\frac{(lm-3)(lm-4)}{12}\right\rceil.$
Proof.
By Theorem 4.3 we have $\mathop{\mathcal{E}_{c}^{-}}(X)\cong K_{lm\cdot n}$. Hence, the result follows from Theorem 5.1, putting $a=lm$ and $b=n$. ∎
Corollary 5.5.
Let $G$ be a finite non-Engel group such that $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m}$. Let $H$ be a finite Engel group with $|H|=l$ and $X=H\times G$. Then $\gamma(\mathop{\mathcal{E}_{c}^{-}}(X))=\left\lceil\frac{(lm-3)(lm-4)}{12}\right\rceil.$
Proof.
Note that $K_{m}\cong K_{m\cdot 1}$. By Theorem 4.3, we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{lm\cdot 1}\cong K_{lm}$. Hence, the result follows from (1). ∎
Corollary 5.6.
Let $H$ be a finite Engel group. Let $t$ and $m$ be positive integers with $t\geq 1$, $m\geq 3$ and $m$ is odd. Then
$\mathop{\mathcal{E}_{c}^{-}}(H\times D_{2^{t+1}m)}\cong\mathop{\mathcal{E}_{c}^{-}}(H\times Q_{2^{t+1}m}).$
Let $|H|=l$ and $K=D_{2^{t+1}m}$ or $Q_{2^{t+1}m}$. Then $\mathop{\mathcal{E}_{c}^{-}}(H\times K)\cong K_{lm\cdot 2^{t}}$ and $\mathop{\mathcal{E}_{c}^{-}}(H\times D_{2m})\cong K_{lm}$. Further,
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(H\times K))=\frac{lm(lm-1)(2^{t-1}-1)^{2}}{2}+\left\lceil\frac{(lm-3)(lm-4)}{12}\right\rceil$
and $\gamma(\mathop{\mathcal{E}_{c}^{-}}(H\times D_{2m}))=\left\lceil\frac{(lm-3)(lm-4)}{12}\right\rceil$.
Proof.
By Theorem 4.1 we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m\cdot 2^{t}}$ if $G=D_{2^{t+1}m}$ or $Q_{2^{t+1}m}$, and $\mathop{\mathcal{E}_{c}^{-}}(D_{2m})\cong K_{m}$. In view of this, it is easy to see that
$\mathop{\mathcal{E}_{c}^{-}}(H\times D_{2^{t+1}m})\cong\mathop{\mathcal{E}_{c}^{-}}(H\times Q_{2^{t+1}m}).$
If $|H|=l$ and $K=D_{2^{t+1}m}$ or $Q_{2^{t+1}m}$ then, by Theorem 4.3, $\mathop{\mathcal{E}_{c}^{-}}(H\times K)\cong K_{(}lm\cdot 2^{t})$ and $\mathop{\mathcal{E}_{c}^{-}}(H\times D_{2m})\cong K_{lm}$. By Theorem 5.1 we have
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(H\times K))=\frac{lm(lm-1)}{2}\left\lceil\frac{(2^{t}-2)^{2}}{4}\right\rceil+\left\lceil\frac{(lm-3)(lm-4)}{12}\right\rceil.$
Hence we get the required expression for $\gamma(\mathop{\mathcal{E}_{c}^{-}}(H\times K))$. The expression for $\gamma(\mathop{\mathcal{E}_{c}^{-}}(H\times D_{2m}))$ can be obtained from (1). ∎
Corollary 5.7.
Let $F_{p,q}=\langle a,b\mid a^{p}=b^{q}=1,a^{-1}ba=b^{r}\rangle$ be a group of order $pq$, where $p$ and $q$ are primes such that $q\equiv 1\mod p$ and $r^{p}\equiv 1\mod q$. Let $H$ be a finite Engel group with $|H|=l$. Then $\mathop{\mathcal{E}_{c}^{-}}(H\times F_{p,q})\cong K_{lq\cdot(p-1)}$. Further,
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(H\times F_{p,q}))=\frac{lq(lq-1)}{2}\left\lceil\frac{(p-3)^{2}}{4}\right\rceil+\left\lceil\frac{(lq-3)(lq-4)}{12}\right\rceil.$
Proof.
By Theorem 4.2 we have $\mathop{\mathcal{E}_{c}^{-}}(F_{p,q})=K_{q\cdot(p-1)}$. In view of this and Theorem 4.3, it follows that $\mathop{\mathcal{E}_{c}^{-}}(H\times F_{p,q})\cong K_{lq\cdot(p-1)}$.
Putting $a=lq$ and b=p1b=p-1 in Theorem 5.1, we have
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(H\times F_{p,q}))=\frac{lq(lq-1)}{2}\left\lceil\frac{((p-1)-2)^{2}}{4}\right\rceil+\left\lceil\frac{(lq-3)(lq-4)}{12}\right\rceil.$
Hence, we get the required expression for $\gamma(\mathop{\mathcal{E}_{c}^{-}}(H\times F_{p,q}))$. ∎
The remaining part of this section is devoted to the characterization of finite groups through the genus of $\mathop{\mathcal{E}_{c}^{-}}(G)$. We first recall the following characterization of finite groups obtained by Abdollahi [2].
Theorem 5.8 ([2], Theorem 3.1).
Let $G$ be a finite non-Engel group. Then $\mathop{\mathcal{E}_{c}^{-}}(G)$ is planar if and only if $G\cong D_{6},D_{12}$ or $Q_{12}$.
In the next two results we characterize the groups $G=D_{2^{t+1}m}$, $Q_{2^{t+1}m}$, $D_{2m}$ (where $t\geq 1$, $m\geq 3$ and $m$ is odd) or $F_{p,q}$ such that the graphs $\mathop{\mathcal{E}_{c}^{-}}(G)$ are planar, toroidal, double-toroidal and triple-toroidal.
Theorem 5.9.
If $G$ is isomorphic to the group $D_{2m}$, where $m\geq 3$ and $m$ is odd, then
  1. (a)
    $\mathop{\mathcal{E}_{c}^{-}}(G)$ is planar if and only if $m=3$.
  2. (b)
    $\mathop{\mathcal{E}_{c}^{-}}(G)$ is toroidal if and only if $m=5,7$.
  3. (c)
    $\mathop{\mathcal{E}_{c}^{-}}(G)$ is not double-toroidal.
  4. (d)
    $\mathop{\mathcal{E}_{c}^{-}}(G)$ is triple-toroidal if and only if $m=9$.
  5. (e)
    $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\geq 5$ for $m\geq 11$.
Proof.
If $G$ is isomorphic to $D_{2m}$, $m\geq 3$ and $m$ is odd, then by Theorem 5.2 we have $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\lceil\frac{(m-3)(m-4)}{12}\right\rceil$. If $m=3$ then we have $\frac{(m-3)(m-4)}{12}=0$. Therefore, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=0$. If $m=5$ and $7$ then $\frac{(m-3)(m-4)}{12}=\frac{1}{6}$ and $1$ respectively. Therefore, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=1$. If $m=9$ then $\frac{(m-3)(m-4)}{12}=2.5$. Therefore, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=3$. If $m\geq 11$ then $\frac{(m-3)(m-4)}{12}=4+\frac{2}{3}$. Therefore, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\geq 5$. Hence, the result follows. ∎
Theorem 5.10.
If $G$ is isomorphic to the group $D_{2^{t+1}m}$ or $Q_{2^{t+1}m}$, where $t\geq 1$ and $m\geq 3$ is odd, then
  1. (a)
    $\mathop{\mathcal{E}_{c}^{-}}(G)$ is planar if and only if $t=1$ and $m=3$.
  2. (b)
    $\mathop{\mathcal{E}_{c}^{-}}(G)$ is toroidal if and only if $t=1$ and $m=5,7$.
  3. (c)
    $\mathop{\mathcal{E}_{c}^{-}}(G)$ is not double-toroidal.
  4. (d)
    $\mathop{\mathcal{E}_{c}^{-}}(G)$ is triple-toroidal if and only if $t=1$, $m=9$ and $t=2$, $m=3$.
  5. (e)
    $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\geq 5$ for any other values of $t$ and $m$.
Proof.
If $G$ is isomorphic to $D_{2^{t+1}m}$ or $Q_{2^{t+1}m}$, where $t\geq 1$, $m\geq 3$ and $m$ is odd, then by Theorem 5.2 we have
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=\frac{m(m-1)(2^{t-1}-1)^{2}}{2}+\left\lceil\frac{(m-3)(m-4)}{12}\right\rceil.$
We consider the following cases.
Case 1. $t=1$
In this case we have $(2^{t-1}-1)^{2}=0$ and so $\frac{m(m-1)(2^{t-1}-1)^{2}}{2}=0$. If $m=3$ then (m3)(m4)=0(m-3)(m-4)=0. Therefore, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=0$. If $m=5,7$ then $\left\lceil\frac{(m-3)(m-4)}{12}\right\rceil=1$. Therefore, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=1$. If $m=9$ then $\left\lceil\frac{(m-3)(m-4)}{12}\right\rceil=3$. Therefore, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=3$. If $m\geq 11$ then $\left\lceil\frac{(m-3)(m-4)}{12}\right\rceil\geq 5$ and so $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\geq 5$.
Case 2. $t=2$
In this case we have $(2^{t-1}-1)^{2}=1$. If $m=3$ then $\frac{m(m-1)(2^{t-1}-1)^{2}}{2}=3$ and (m3)(m4)=0(m-3)(m-4)=0. Therefore, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=3$. If $m\geq 5$ then $\frac{m(m-1)(2^{t-1}-1)^{2}}{2}\geq 10$ and $\left\lceil\frac{(m-3)(m-4)}{12}\right\rceil\geq 1$. Therefore, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\geq 11$.
Case 3. $t\geq 3$
In this case $(2^{t-1}-1)^{2}\geq 9$ and $m(m-1)\geq 6$. Therefore, $\frac{m(m-1)(2^{t-1}-1)^{2}}{2}\geq 27$. Hence, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\geq 27$ since $\left\lceil\frac{(m-3)(m-4)}{12}\right\rceil\geq 0$. Thus we get the required result. ∎
Theorem 5.11.
If $G$ is isomorphic to the group $F_{p,q}$ then
  1. (a)
    $\mathop{\mathcal{E}_{c}^{-}}(G)$ is planar if and only if $p=2$ and $q=3$.
  2. (b)
    $\mathop{\mathcal{E}_{c}^{-}}(G)$ is toroidal if and only if $p=2$ and $q=5,7$; $p=3$ and $q=7$.
  3. (c)
    $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\geq 5$ for any other values of $p$ and $q$. Thus, $\mathop{\mathcal{E}_{c}^{-}}(G)$ is neither double-toroidal nor triple-toroidal.
Proof.
We consider the following cases.
Case 1. $p=2$
By Theorem 5.3 we have $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\lceil\frac{(q-3)(q-4)}{12}\right\rceil$. If $q=3$ then $\frac{(q-3)(q-4)}{12}=0$, Therefore, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=0$. If $q=5$ and $7$ then $\frac{(q-3)(q-4)}{12}=\frac{1}{6}$ and $1$ respectively. Therefore, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=1$. If $q\geq 11$ then $\frac{(q-3)(q-4)}{12}\geq 4+\frac{2}{3}$. Therefore, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\geq 5$.
Case 2. $p=3$
By Theorem 5.3 we have
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=\frac{q(q-1)}{2}\left\lceil\frac{(p-3)^{2}}{4}\right\rceil+\left\lceil\frac{(q-3)(q-4)}{12}\right\rceil.$
In this case we have $\frac{q(q-1)}{2}\left\lceil\frac{(p-3)^{2}}{4}\right\rceil=0$. If $q=7$ then $\frac{(q-3)(q-4)}{12}=1$. Therefore, $\left\lceil\frac{(q-3)(q-4)}{12}\right\rceil=1$ and so $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=1$. If $q\geq 13$ then $\frac{(q-3)(q-4)}{12}\geq 7.5$. Therefore, $\left\lceil\frac{(q-3)(q-4)}{12}\right\rceil\geq 8$ and so $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\geq 8$.
Case 3. $p\geq 5$
By Theorem 5.3 we have
$\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))=\frac{q(q-1)}{2}\left\lceil\frac{(p-3)^{2}}{4}\right\rceil+\left\lceil\frac{(q-3)(q-4)}{12}\right\rceil.$
In this case we have $\left\lceil\frac{(p-3)^{2}}{4}\right\rceil\geq 1$. If $q\geq 11$ then $\frac{q(q-1)}{2}\geq 55$ and $\left\lceil\frac{(q-3)(q-4)}{12}\right\rceil\geq 5$. Therefore, $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\geq 60$. Hence, the result follows. ∎
Theorem 5.12.
Let $G$ be a finite non-Engel group such that $\omega(\mathop{\mathcal{E}_{c}^{-}}(G))\leq 4$. Then
  1. (a)
    $\mathop{\mathcal{E}_{c}^{-}}(G)$ is toroidal if and only if $G\cong A_{4},C_{3}\times D_{6}$,
  2. (b)
    $\mathop{\mathcal{E}_{c}^{-}}(G)$ is not double-toroidal.
Proof.
Since $\omega(\mathop{\mathcal{E}_{c}^{-}}(G))\leq 4$, by [1, Proposition 1.4] and [2, Theorem 1.2], we have $\frac{G}{Z^{*}(G)}\cong D_{6}$ or $A_{4}$.
Suppose $\frac{G}{Z^{*}(G)}\cong D_{6}$. Let $\bar{G}=\frac{G}{Z^{*}(G)}$ and $\bar{x}=xZ^{*}(G)$ for every $x\in G$. Since $\bar{G}\cong D_{6}$, we have $\mathop{\mathcal{E}_{c}^{-}}(\bar{G})\cong\mathop{\mathcal{E}_{c}^{-}}(D_{6})$. Let $\phi:\mathop{\mathcal{E}_{c}^{-}}(\bar{G})\longrightarrow\mathop{\mathcal{E}_{c}^{-}}(D_{6})$ be a graph isomorphism. Let $a,b\in D_{6}$ such that $D_{6}=\langle a,b\rangle$ and $a^{3}=b^{2}=1$. Then $\{b,ab,a^{2}b\}$ forms a clique in $\mathop{\mathcal{E}_{c}^{-}}(D_{6})$. Therefore $\{\bar{x}_{1}=\phi^{-1}(b),\bar{x}_{2}=\phi^{-1}(ab),\bar{x}_{3}=\phi^{-1}(a^{2}b)\}$ forms a clique of $\mathop{\mathcal{E}_{c}^{-}}(\bar{G})$. By Theorem 3.2, every elements of $x_{i}Z^{*}(G)$ is adjacent to every elements of $x_{j}Z^{*}(G)$, for $i,j\in\{0,1,2\},i\neq j$. Thus $K_{m,m,m}$ is a subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$, where $m=|Z^{*}(G)|$. Therefore, by (3), we have
$\gamma(K_{m,m,m})=\frac{(m-1)(m-2)}{2}\leq\gamma(\mathop{\mathcal{E}_{c}^{-}}(G)).$ (4)
Suppose that $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\leq 2$. In that case, by (4), we get $m\leq 3$.
If $m=1$, then $G$ is a non-nilpotent group of order $6$. Thus $G\cong D_{6}$ and $\mathop{\mathcal{E}_{c}^{-}}(D_{6})$ is planar.
If $m=2$, then $G$ is a non-nilpotent group of order $12$ with $|Z^{*}(G)|=2$. Thus $G\cong D_{12},Q_{12}$. Therefore, $\mathcal{E}_{G}[G\setminus L(G)]$ is planar
If $m=3$, then $G$ is a non-nilpotent graph of order $18$ with $|Z^{*}(G)|=3$. Thus $G\cong C_{3}\times D_{6}$. Here $\mathop{\mathcal{E}_{c}^{-}}(C_{3}\times D_{6})\cong K_{3,3,3}$ and $\gamma(K_{3,3,3})=\frac{(3-1)(3-2)}{2}=1$. Thus $\mathop{\mathcal{E}_{c}^{-}}(C_{3}\times D_{6})$ is toroidal. An embedding of $\mathop{\mathcal{E}_{c}^{-}}(C_{3}\times D_{6})$ on a torus is shown in Figure 1, where $C_{3}=\langle a\rangle$.
[…]
Figure 1. Embedding of $\mathop{\mathcal{E}_{c}^{-}}(C_{3}\times D_{6})$ on a torus.
(2,3,4)(2,3,4)(1,2,3)(1,2,3)(1,3,4)(1,3,4)(1,3,2)(1,3,2)(1,4,3)(1,4,3)(1,4,2)(1,4,2)(2,4,3)(2,4,3)$(1,2,4)$
Figure 2. $\mathop{\mathcal{E}_{c}^{-}}(A_{4})$
(2,3,4)(2,3,4)(2,3,4)(2,3,4)(1,2,3)(1,2,3)(1,2,3)(1,2,3)(1,3,4)(1,3,4)(1,3,4)(1,3,4)(1,3,2)(1,3,2)(1,3,2)(1,3,2)(1,4,2)(1,4,2)(1,2,4)(1,2,4)(2,4,3)(2,4,3)(1,4,3)(1,4,3)aaaabbbbccccddddeeee
Figure 3. Embedding of $\mathop{\mathcal{E}_{c}^{-}}(A_{4})$ on a torus.
Suppose $\frac{G}{Z^{*}(G)}\cong A_{4}$. Suppose that $|Z^{*}(G)|\geq 2$. Let $\bar{G}=\frac{G}{Z^{*}(G)}$ and $\bar{x}=xZ^{*}(G)$ for every $x\in G$. Since $\bar{G}\cong A_{4}$, we have $\mathop{\mathcal{E}_{c}^{-}}(\bar{G})\cong\mathop{\mathcal{E}_{c}^{-}}(A_{4})$. Let $\phi:\mathop{\mathcal{E}_{c}^{-}}(\bar{G})\longrightarrow\mathop{\mathcal{E}_{c}^{-}}(A_{4})$ be a graph isomorphism. As we see in Figure 2, $\mathop{\mathcal{E}_{c}^{-}}(A_{4})$ has a subgraph isomorphic to $K_{4,4}$ with parts $H=\{(2,3,4),(1,2,4),(2,4,3),(1,4,2)\}$ and $K=\{(1,2,3),(1,3,4),(1,3,2),(1,4,3)\}$. Therefore, $\mathop{\mathcal{E}_{c}^{-}}(\bar{G})$ has a subgraph isomorphic to $K_{4,4}$ with parts
$\bar{H}=\{\phi^{-1}(2,3,4),\phi^{-1}(1,2,4),\phi^{-1}(2,4,3),\phi^{-1}(1,4,2)\}$
and
$\bar{K}=\{\phi^{-1}(1,2,3),\phi^{-1}(1,3,4),\phi^{-1}(1,3,2),\phi^{-1}(1,4,3)\}.$
Let $z\in Z^{*}(G),z\neq 1$. Then, by Theorem 3.2, every elements of $\bar{H}\cup z\bar{H}$ is adjacent to every elements of $\bar{K}\cup z\bar{K}$, showing that $K_{8,8}$ is a subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$.
Thus $\gamma(\mathop{\mathcal{E}_{c}^{-}}(G))\geq\gamma(K_{8,8})=9$, a contradiction. Thus $|Z^{*}(G)|=1$ and so $G\cong A_{4}$. An embedding of the graph $\mathop{\mathcal{E}_{c}^{-}}(A_{4})$ on a torus is shown in Figure 3. This completes the proof. ∎
Let $N_{k}$ denote the surface formed by a connected sum of $k$ projective planes ($k\geq 1$). The number $k$ is called the crosscap of $N_{k}$. A simple graph $\Gamma$ which can be embedded in $N_{k}$ but not in $N_{k-1}$, is call a graph of crosscap $\bar{\gamma}(\Gamma)=k$. A graph $\Gamma$ with $\bar{\gamma}(\Gamma)=1$ is called a projective graph. The following theorem is useful in the next result.
Theorem 5.13.
([6, 17]). For positive integers $m$ and $n$, we have
  1. (a)
    $\bar{\gamma}(K_{m,n})=\lceil\frac{(m-2)(n-2)}{2}\rceil$ if $m,n\geq 2$.
  2. (b)
[…]
We conclude this section by determining all finite non-Engel groups $G$ (up to isomorphism) such that $\omega(\mathop{\mathcal{E}_{c}^{-}}(G))\leq 4$ and $\mathop{\mathcal{E}_{c}^{-}}(G)$ is projective.
Theorem 5.14.
Let $G$ be a finite non-Engel group such that $\omega(\mathop{\mathcal{E}_{c}^{-}}(G))\leq 4$. Then $\mathop{\mathcal{E}_{c}^{-}}(G)$ is projective if and only if $G\cong D_{6},D_{12},Q_{12}$.
Proof.
By [1, Proposition 1.4] and [2, Theorem 1.2], we have $G/Z^{*}(G)\cong D_{6}$ or $A_{4}$. Suppose $G/Z^{*}(G)\cong A_{4}$. As seen in Figure 2, $\mathop{\mathcal{E}_{c}^{-}}(A_{4})$ has a subgraph isomorphic to $K_{4,4}$. Thus $\mathop{\mathcal{E}_{c}^{-}}(G/Z^{*}(G))$ has a subgraph isomorphic to $K_{4,4}$. Note that $|v(\mathop{\mathcal{E}_{c}^{-}}(G/Z^{*}(G)))|=|v(\mathop{\mathcal{E}_{c}^{-}}(A_{4}))|=8$. Let $v(\mathop{\mathcal{E}_{c}^{-}}(G/Z^{*}(G)))=\{\bar{a_{1}},\bar{a_{2}},\dots,\bar{a_{8}}\}$. Then the induced subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$ by the set $\{a_{1},a_{2},\dots,a_{8}\}$ has a subgraph isomorphic to $K_{4,4}$. Thus $\bar{\gamma}(\mathop{\mathcal{E}_{c}^{-}}(G))\geq\bar{\gamma}(K_{4,4})=2$, a contradiction. Thus $G/Z^{*}(G)\cong D_{6}$.
Let $m=|Z^{*}(G)|$. As seen in the proof of Theorem 5.12, $K_{m,m,m}$ is a subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$. Thus if $m\geq 3$, then $K_{6,3}$ is a subgraph of $\mathop{\mathcal{E}_{c}^{-}}(G)$. But $\bar{\gamma}(K_{6,3})=2$. Therefore, $m\leq 2$. If $m=1$, then $G\cong D_{6}$ and if $m=2$, then $G\cong D_{12},Q_{12}$. Note that $\mathop{\mathcal{E}_{c}^{-}}(D_{6})\cong K_{3}$ is projective. Also, from Theorem 4.1, $\mathop{\mathcal{E}_{c}^{-}}(D_{12})\cong\mathop{\mathcal{E}_{c}^{-}}(Q_{12})\cong K_{3\cdot 2}$ which is projective. This completes the proof. ∎
6. Various spectra and energies of $\mathop{\mathcal{E}_{c}^{-}}(G)$
Let $A(\Gamma)$ and $D(\Gamma)$ be the adjacency matrix and degree matrix of a graph $\Gamma$. Let $L(\Gamma):=D(\Gamma)-A(\Gamma)$ and $Q(\Gamma):=D(\Gamma)+A(\Gamma)$ be the Laplacian and signless Laplacian matrices of $\Gamma$. The characteristic polynomials of $A(\Gamma),L(\Gamma)$ and $Q(\Gamma)$ are called characteristic polynomial (denoted by $P(\Gamma,x)$), Laplacian polynomial (denoted by $P_{L}(\Gamma,x)$) and signless Laplacian polynomial (denoted by $P_{Q}(\Gamma,x)$) of $\Gamma$ respectively. The set of all the roots (with multiplicities) of $P(\Gamma,x)$, $P_{L}(\Gamma,x)$ and $P_{Q}(\Gamma,x)$ are called spectrum, Laplacian spectrum and signless Laplacian spectrum of $\Gamma$ denoted by $\operatorname{Spec}(\Gamma),\operatorname{L-spec}-Spec(\Gamma)$ and $\operatorname{Q-spec}-Spec(\Gamma)$ respectively. If $\{(\alpha_{1})^{n_{1}},(\alpha_{2})^{n_{2}},\dots,(\alpha_{k})^{n_{k}}\}$ represents $\operatorname{Spec}(\Gamma),\operatorname{L-spec}-Spec(\Gamma)$ or $\operatorname{Q-spec}-Spec(\Gamma)$ then $\alpha_{i}$’s are roots of $P(\Gamma,x)$, $P_{L}(\Gamma,x)$ or $P_{Q}(\Gamma,x)$ with multiplicities $n_{i}$. The energy, Laplacian energy and signless Laplacian energy of $\Gamma$ denoted by $E(\Gamma)$, $LE(\Gamma)$ and $LE^{+}(\Gamma)$ respectively are defined as
$E(\Gamma)=\sum_{\lambda\in\operatorname{Spec}(\Gamma)}|\lambda|,$
$LE(\Gamma)=\sum_{\mu\in\operatorname{L-spec}-Spec(\Gamma)}\left|\lambda-\frac{2|e(\Gamma)|}{|v(\Gamma)|}\right|\quad\text{and}\quad LE^{+}(\Gamma)=\sum_{\nu\in\operatorname{Q-spec}-Spec(\Gamma)}\left|\mu-\frac{2|e(\Gamma)|}{|v(\Gamma)|}\right|.$
A graph is called hyperenergetic or hypoenergetic if $E(\Gamma)>E(K_{|v(\Gamma)|})$ or $E(\Gamma)<|v(\Gamma)|$ respectively. It is not known whether co-Engel graphs of finite groups are hyperenergetic or hypoenergetic in general. However, we shall show that the $\mathop{\mathcal{E}_{c}^{-}}(G)$ for finite groups considered in this paper are neither hyperenergetic nor hypoenergetic. In [12], Gutman et al. conjectured that $E(\Gamma)\leq LE(\Gamma)$ for any finite graph $\Gamma$ (which is known as the E-LE conjecture). In general, this conjecture is false. It is not known whether co-Engel graphs of finite groups satisfy the E-LE conjecture. Our computations show that the co-Engel graphs of finite groups considered in this paper satisfy this conjecture.
The following two well-known results are useful in computing various spectra and energies of $\mathop{\mathcal{E}_{c}^{-}}(G)$ for the groups considered in Section 4.
Lemma 6.1.
If $\Gamma=K_{n}$ then
$P(\Gamma,x)=(x+1)^{n-1}\left(x-(n-1)\right),\quad P_{L}(\Gamma,x)=x(x-n)^{n-1}\text{ and }$
$P_{Q}(\Gamma,x)=(x-(n-2))^{n-1}(x-2(n-1)).$
Lemma 6.2.
If $\Gamma=K_{a\cdot b}$ then
$P(\Gamma,x)=x^{a(b-1)}(x+b)^{a-1}\left(x-b(a-1)\right),$
$P_{L}(\Gamma,x)=x(x-b(a-1))^{a(b-1)}(x-ab)^{a-1}\text{ and }$
$P_{Q}(\Gamma,x)=(x-b(a-1))^{a(b-1)}(x-b(a-2))^{a-1}(x-2b(a-1)).$
If $m\geq 3$ is odd and $G\cong D_{2m}$ then, by Theorem 5.6, we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m}$. Therefore, we get the following result.
Theorem 6.3.
If $G$ is isomorphic to $D_{2m}$, where $m\geq 3$ is odd, then
$\operatorname{Spec}(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(-1)^{m-1},(m-1)^{1}\right\},\quad\operatorname{L-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(0)^{1},m^{m-1}\right\},$
$\operatorname{Q-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))\!=\!\left\{(m-2)^{m-1},(2(m-1))^{1}\right\}\!\text{ and }$
$E(\mathop{\mathcal{E}_{c}^{-}}(G))\!=\!LE(\mathop{\mathcal{E}_{c}^{-}}(G))\!=\!LE^{+}(\mathop{\mathcal{E}_{c}^{-}}(G))=2(m-1).$
Theorem 6.4.
If $G$ is isomorphic to $D_{2^{t+1}m}$ or $Q_{2^{t+1}m}$, where $t\geq 1$ and $m\geq 3$ is odd, then
$\operatorname{Spec}(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(0)^{m(2^{t}-1)},(-2^{t})^{m-1},\left(2^{t}(m-1)\right)^{1}\right\},$
$\operatorname{L-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(0)^{1},(2^{t}(m-1))^{m(2^{t}-1)},(2^{t}m)^{m-1}\right\},$
$\operatorname{Q-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(2^{t}(m-1))^{m(2^{t}-1)},(2^{t}(m-2))^{m-1},(2^{t+1}(m-1))^{1}\right\}$
and $E(\mathop{\mathcal{E}_{c}^{-}}(G))=LE(\mathop{\mathcal{E}_{c}^{-}}(G))=LE^{+}(\mathop{\mathcal{E}_{c}^{-}}(G))=2^{t+1}(m-1)$.
Proof.
By Theorem 5.6 we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m\cdot 2^{t}}$. Using Lemma 6.2 we get
$P(\mathop{\mathcal{E}_{c}^{-}}(G),x)=x^{m(2^{t}-1)}(x+2^{t})^{m-1}\left(x-2^{t}(m-1)\right),$
$P_{L}(\mathop{\mathcal{E}_{c}^{-}}(G),x)=x(x-2^{t}(m-1))^{m(2^{t}-1)}(x-2^{t}m)^{m-1}\text{ and }$
$P_{Q}(\mathop{\mathcal{E}_{c}^{-}}(G),x)=(x-2^{t}(m-1))^{m(2^{t}-1)}(x-2^{t}(m-2))^{m-1}(x-2^{t+1}(m-1)).$
Therefore, $\operatorname{Spec}(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(0)^{m(2^{t}-1)},(-2^{t})^{m-1},\left(2^{t}(m-1)\right)^{1}\right\}$ and so
$E(\mathop{\mathcal{E}_{c}^{-}}(G))=m(2^{t}-1)\times 0+(m-1)2^{t}+2^{t}(m-1)=2^{t+1}(m-1).$
Also, $\operatorname{L-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(0)^{1},(2^{t}(m-1))^{m(2^{t}-1)},(2^{t}m)^{m-1}\right\}$ and
$\operatorname{Q-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(2^{t}(m-1))^{m(2^{t}-1)},(2^{t}(m-2))^{m-1},(2^{t+1}(m-1))^{1}\right\}$.
We have ${|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}=\frac{2^{t}m(2^{t}m-1)}{2}-\frac{2^{t}m(2^{t}-1)}{2}=\frac{2^{t}m(2^{t}m-2^{t})}{2}$ and so $\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}$ $2^{t}(m-1)$. Therefore,
$\left|0-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|=\left|-2^{t}(m-1)\right|=2^{t}(m-1),$
$\left|2^{t}(m-1)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|$ $=\left|2^{t}(m-1)-2^{t}(m-1)\right|=0,$
and
$\left|2^{t}m-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|$ $=\left|2^{t}m-2^{t}(m-1)\right|=2^{t}.$
Therefore,
$LE(\mathop{\mathcal{E}_{c}^{-}}(G))$ $=1\times 2^{t}(m-1)+m(2^{t}-1)\times 0+(m-1)\times 2^{t}$
$=2^{t+1}(m-1).$
Also,
$\left|2^{t}(m-2)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|$ $=\left|2^{t}(m-2)-2^{t}(m-1)\right|=2^{t},$
and
$\left|2^{t+1}(m-1)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|$ $=\left|2^{t+1}(m-1)-2^{t}(m-1)\right|$
$=2^{t}(m-1).$
Therefore
$LE^{+}(\mathop{\mathcal{E}_{c}^{-}}(G))$ $=m(2^{t}-1)\times 0+(m-1)\times 2^{t}+1\times 2^{t}(m-1)$
$=2^{t+1}(m-1).$
Theorem 6.5.
If $G$ is isomorphic to $F_{p,q}$, then
$\operatorname{Spec}(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(0)^{q(p-2)},(-(p-1))^{q-1},\left((p-1)(q-1)\right)^{1}\right\},$
$\operatorname{L-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(0)^{1},((p-1)(q-1))^{q(p-2)},(q(p-1))^{q-1}\right\},$
$\operatorname{Q-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))\!=\!\left\{\!((p\!-\!1)(q-1))^{q(p-2)},((p-1)(q-2))^{q-1},(2(p-1)(q-1))^{1}\!\right\},$
 $\text{ and }E(\mathop{\mathcal{E}_{c}^{-}}(G))=LE(\mathop{\mathcal{E}_{c}^{-}}(G))=LE^{+}(\mathop{\mathcal{E}_{c}^{-}}(G))=2(p-1)(q-1).$
Proof.
By Theorem 4.2 we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{q\cdot(p-1)}$. Using Lemma 6.2 we get
$P(\mathop{\mathcal{E}_{c}^{-}}(G),x)=x^{q(p-2)}(x+(p-1))^{q-1}\left(x-(p-1)(q-1)\right),$
$P_{L}(\mathop{\mathcal{E}_{c}^{-}}(G),x)=x(x-(p-1)(q-1))^{q(p-2)}(x-q(p-1))^{q-1}\text{ and }$
$P_{Q}(\mathop{\mathcal{E}_{c}^{-}}(G),x)=(x-(p-1)(q-1))^{q(p-2)}(x-(p-1)(q-2))^{q-1}(x-2(p-1)(q-1)).$
Therefore, $\operatorname{Spec}(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(0)^{q(p-2)},(-(p-1))^{q-1},\left((p-1)(q-1)\right)^{1}\right\}$ and so
$E(\mathop{\mathcal{E}_{c}^{-}}(G))=q(p-2)\times 0+(q-1)(p-1)+(p-1)(q-1)=2(p-1)(q-1).$
Also, $\operatorname{L-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{(0)^{1},((p-1)(q-1))^{q(p-2)},(q(p-1))^{q-1}\right\}$ and
$\operatorname{Q-spec}-Spec(\mathop{\mathcal{E}_{c}^{-}}(G))=\left\{((p-1)(q-1))^{q(p-2)},((p-1)(q-2))^{q-1},(2(p-1)(q-1))^{1}\right\}$.
We have ${|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}=\frac{q(p-1)^{2}(q-1)}{2}$ and so $\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}=(p-1)(q-1)$. Therefore,
$\left|0-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|=(p-1)(q-1),$
$\left|(p-1)(q-1)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|=\left|(p-1)(q-1)-(p-1)(q-1)\right|=0,$
and
$\left|q(p-1)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|=\left|q(p-1)-(p-1)(q-1)\right|=(p-1).$
Therefore,
$LE(\mathop{\mathcal{E}_{c}^{-}}(G))=1\times(p-1)(q-1)+q(p-2)\times 0+(q-1)\times(p-1)=2(p-1)(q-1).$
Also,
$\left|(p-1)(q-2)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|$ $=\left|(p-1)(q-2)-(p-1)(q-1)\right|$
$=(p-1),$
and
$\left|2(p-1)(q-1)-\frac{2|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}\right|$ $=\left|2(p-1)(q-1)-(p-1)(q-1)\right|$
$=(p-1)(q-1).$
and hence
$LE^{+}(\mathop{\mathcal{E}_{c}^{-}}(G))$ $=q(p-2)\times 0+(q-1)\times(p-1)+1\times(p-1)(q-1)$
$=2(p-1)(q-1).$
Corollary 6.6.
If $G$ is isomorphic to $D_{2^{t+1}m}$ or $Q_{2^{t+1}m}$, where $t\geq 1$ and $m\geq 3$ is odd, then $\mathop{\mathcal{E}_{c}^{-}}(G)$ is neither hyperenergetic nor hypoenergetic.
Proof.
We have $|v(\mathop{\mathcal{E}_{c}^{-}}(G))|=2^{t}m$ and
$E(K_{2^{t}m})=2(2^{t}m-1).$
By Theorem 6.4 we get
$E(\mathop{\mathcal{E}_{c}^{-}}(G))-E(K_{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|})$ $=2^{t+1}(m-1)-2(m2^{t}-1)$
$=-2(2^{t}-1)<0.$
Therefore, $\mathop{\mathcal{E}_{c}^{-}}(G)$ is not hyperenergetic. We also have
$E(\mathop{\mathcal{E}_{c}^{-}}(G))-|v(\mathop{\mathcal{E}_{c}^{-}}(G))|$ $=2^{t+1}(m-1)-2^{t}m$
$=2^{t}(m-2)>0.$
Hence, $\mathop{\mathcal{E}_{c}^{-}}(G)$ is not hypoenergetic. ∎
Corollary 6.7.
If $G$ is isomorphic to $F_{p,q}$, then $\mathop{\mathcal{E}_{c}^{-}}(G)$ is neither hyperenergetic nor hypoenergetic.
Proof.
We have $|v(\mathop{\mathcal{E}_{c}^{-}}(G))|=q(p-1)$ and
$E(K_{q(p-1)})=2(q(p-1)-1).$
By Theorem 6.5 we get
$E(\mathop{\mathcal{E}_{c}^{-}}(G))-E(K_{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|})$ $=2(p-1)(q-1)-2(q(p-1)-1)$
$=-2(p-2)\leq 0.$
Therefore, $\mathop{\mathcal{E}_{c}^{-}}(G)$ is not hyperenergetic. We also have
$E(\mathop{\mathcal{E}_{c}^{-}}(G))-|v(\mathop{\mathcal{E}_{c}^{-}}(G))|$ $=2(p-1)(q-1)-q(p-1)$
$=(p-1)(q-2)>0.$
Hence, $\mathop{\mathcal{E}_{c}^{-}}(G)$ is not hypoenergetic. ∎
7. Zagreb indices of $\mathop{\mathcal{E}_{c}^{-}}(G)$
In this section, we compute Zagreb indices of $\mathop{\mathcal{E}_{c}^{-}}(G)$ for certain dihedral groups, dicyclic groups and the group $F_{p,q}$ and check whether they satisfy the Hansen–Vukičević conjecture. The first and second Zagreb indices of a simple undirected graph $\Gamma$, denoted by $M_{1}(\Gamma)$ and $M_{2}(\Gamma)$ respectively, are given by
$M_{1}(\Gamma)=\sum\limits_{v\in v(\Gamma)}\deg(v)^{2}\text{ and }M_{2}(\Gamma)=\sum\limits_{uv\in e(\Gamma)}\deg(u)\deg(v),$
where $\deg(v)$ is the degree of $v$. The following lemma is useful in obtaining our results.
Lemma 7.1.
If $\Gamma$ is isomorphic to the graph $K_{a\cdot b}$ then $M_{1}(\Gamma)=a(a-1)^{2}b^{3}$ and $M_{2}(\Gamma)=\frac{a(a-1)^{3}b^{4}}{2}$.
In 2007, Hansen and Vukičević [14] conjectured that
$\dfrac{M_{2}(\Gamma)}{|e(\Gamma)|}\geq\dfrac{M_{1}(\Gamma)}{|v(\Gamma)|}.$ (5)
In general, (5) is not true. However, it is not known whether (5) satisfies for $\mathop{\mathcal{E}_{c}^{-}}(G)$. The following results show that (5) is true for $\mathop{\mathcal{E}_{c}^{-}}(G)$ if $G$ is isomorphic to $D_{2^{t+1}m}$, $Q_{2^{t+1}m}$ and $F_{p,q}$, where $t\geq 1$, $m\geq 3$ is odd and $p$, $q$ are primes such that $q\equiv 1\mod p$.
Theorem 7.2.
If $G$ is isomorphic to $D_{2^{t+1}m}$ or $Q_{2^{t+1}m}$, where $t\geq 1$ and $m\geq 3$ is odd, then
$M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))=2^{3t}m(m-1)^{2}\text{ and }M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))=2^{4t-1}m(m-1)^{3}$
Further, $\frac{M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}=2^{2t}(m-1)^{2}=\frac{M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}$.
Proof.
By Theorem 4.1 we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m\cdot 2^{t}}$. Therefore, by Lemma 7.1, we get
$M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))=2^{3t}m(m-1)^{2}.$
and
$M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))=2^{4t-1}m(m-1)^{3}.$
Here $|v(\mathop{\mathcal{E}_{c}^{-}}(G))|=2^{t}m$ and $|e(\mathop{\mathcal{E}_{c}^{-}}(G))|=2^{2t-1}m(m-1)$. Therefore
$\frac{M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}=\frac{2^{3t}m(m-1)^{2}}{2^{t}m}=2^{2t}(m-1)^{2}.$
and
$\frac{M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}=\frac{2^{4t-1}m(m-1)^{3}}{2^{2t-1}m(m-1}=2^{2t}(m-1)^{2}.$
Hence, the result follows. ∎
Theorem 7.3.
If $G$ is isomorphic to $F_{p,q}$, then
$M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))=q(q-1)^{2}(p-1)^{3}\text{ and }M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))=\frac{q(q-1)^{3}(p-1)^{4}}{2}$
Further, $\frac{M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}=(q-1)^{2}(p-1)^{2}=\frac{M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}$.
Proof.
By Theorem 4.2 we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{q\cdot(p-1)}$. Therefore, by Lemma 7.1, we get
$M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))=q(q-1)^{2}(p-1)^{3}.$
and
$M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))=\frac{q(q-1)^{3}(p-1)^{4}}{2}.$
Here $|v(\mathop{\mathcal{E}_{c}^{-}}(G))|=q(p-1)$ and $|e(\mathop{\mathcal{E}_{c}^{-}}(G))|=\frac{q(p-1)^{2}(q-1)}{2}$. Therefore
$\frac{M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}=\frac{q(q-1)^{2}(p-1)^{3}}{q(p-1)}=(q-1)^{2}(p-1)^{2}.$
and
$\frac{M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}=\frac{q(q-1)^{3}(p-1)^{4}}{q(p-1)^{2}(q-1)}=(q-1)^{2}(p-1)^{2}.$
Hence, the result follows. ∎